# Knowledge Enhanced Conditional Imputation for Healthcare Time-series

### 27 Dec 2023 v1

@@ Abstract
Extensive experimental evaluations demonstrate CSAI’s superior performance over existing methods, particularly in managing the complexities of missing data in healthcare time series. Furthermore, we have redefined the time series interpolation task, ensuring greater alignment with clinical demands. In our commitment to collaborative advancement, the code for CSAI will be made available, facilitating further research and application in the field of healthcare analytics. The code will be made available at https://github.com/LINGLONGQIAN/CSAI.
@@ 5 Experiments
In this section, we evaluate CSAI performance against the state-of-the-art baselines using four real-world healthcare datasets. We compare CSAI against BRITS, BRITS-GRU, GRUD, V-RIN(full) and MRNN. In all experiments, only the best model from each study is used for comparison. Although it would have been interesting to know how $E^{2}$ compares with other models, especially given that [29] does not provide a quantitative comparison with BRITS, we were unable to include $E^{2}$ in our study because the publicly-available code is based on the outdated TensorFlow 1.7 and Python 2.7, which are incompatible with our available GPU hardware.
@@ 5.3 Experimental Results
Table 2 reflects CSAI’s robust classification capabilities, with competitive Area Under the Curve (AUC) scores across various masking ratios. In the mimic_59f dataset, CSAI exhibits superior performance, particularly at a 10% masking ratio. This result aligns with our hypothesis that embedding domain knowledge into the model’s structure enhances its predictive accuracy, a critical aspect of healthcare decision support systems.
Our experimental evaluation reveals that CSAI maintains high performance in both imputation and classification tasks, signifying its adaptability to the intrinsic complexities of healthcare datasets. The consistency in performance with increasing masking ratios indicates the model’s resilience to data sparsity, a common challenge in real-world healthcare settings. The model’s domain knowledge integration and advanced imputation strategies contribute to its robustness, reinforcing the importance of specialized approaches for healthcare analytics.
Despite the overall strong performance of CSAI, certain datasets and masking ratios yielded less optimal results, meriting careful analysis. For instance, in the physionet dataset at a 20% masking ratio, CSAI did not achieve the highest AUC score. This suggests potential areas for model refinement, particularly in handling datasets with substantial missingness. Similarly, in the mimic_89f dataset under the same masking condition, CSAI’s performance lagged slightly behind Brits_GRU. These instances point to the possibility that CSAI could benefit from further tuning of its domain knowledge integration mechanisms, or from additional training data, to enhance its robustness across all levels of data sparsity.
@@
BVRIN 0.8869942305834426(0) 0.8860338953927615(0) 0.8841467199689977(0) 0.8318772129986762(0) 0.8300245958652113(0) 0.8243715327663429(0)
0.8857860601179969(10) 0.8815291839908802(10) 0.8812348535321004(10) 0.8302246518225975(10) 0.8277489166942631(10) 0.8239595531167085(10)
Brits 0.8880503633414845(0) / 0.8862824221267139(0) 0.8347218170342(0) 0.8254274124061122(0) 0.8310903756879437(0)
0.8895208094995246(10) 0.8882202129602428(10) 0.8879479677632803(10) 0.832124270188(10) 0.826299316340254(10) 0.8262743541111952(10)
Brits_GRU 0.8858896666666668(0) 0.8851533333333333(0) 0.8855963333333333(0) 0.8275031457832767(0) 0.8259190600205351(0) 0.8268552412968461(0)
0.8847200000000001(10) 0.8859573333333334(10) 0.8864126666666667(10) 0.8285739607100067(10) 0.8255457714836528(10) 0.8249535939343665(10)
@@ 5.4.2 Adjustment Factor Comparison
The experimental results depicted in the table indicate a nuanced relationship between the adjustment factor in the non-uniform masking strategy and the performance on imputation and classification tasks. An optimal adjustment factor demonstrated to be around 5, results in the lowest imputation error, suggesting that a balanced representation of features is crucial for accuracy. For classification, however, increasing the adjustment factor leads to higher errors and a marginal decline in AUC, highlighting that excessive weighting may not uniformly enhance performance across different machine learning tasks. These findings reveal the delicate interplay between feature representation adjustments and task-specific model efficacy, underscoring the importance of calibration in the non-uniform masking approach for complex time series data.

### 04 Jan 2024 v2

@@ Abstract
Extensive experimental evaluations demonstrate that CSAI outperforms existing methods, particularly in managing the complexities of missing data in healthcare time series. Furthermore, we have redefined the time series interpolation task, ensuring greater alignment with clinical demands. In our commitment to collaborative advancement, the code for CSAI will be made available, facilitating further research and application in the field of healthcare analytics. The code will be made available at https://github.com/LINGLONGQIAN/CSAI.
@@ 4.2 Conditional Knowledge Embedding
Integration of Median Time Gap To enhance the model’s healthcare applicability, we have integrated the median time gap $\boldsymbol{\tau}$ for a specific healthcare feature. This integration leads to a tailored adjustment, giving more weight to observations closer to the median time gap and less to those further away. This adjusted function is formulated as:
@@ 5 Experiments
In this study, we rigorously evaluate the performance of our Conditional Self-Attention Imputation (CSAI) model against a selection of state-of-the-art baselines in the context of four real-world healthcare datasets. Our comparison includes an array of established models: BRITS, BRITS-GRU, GRUD, V-RIN(full), and MRNN. It’s important to note that we specifically include BRITS-GRU as a baseline due to the original BRITS model utilizing LSTM cells, while other baselines employ GRU cells. This addition ensures a fair and comprehensive comparison across similar architectures.
For each experimental setup, we select only the best-performing model from each baseline study for a more focused and meaningful comparison. Although a comparison with $E^{2}$ would have provided additional insights, especially since it did not quantitatively compare it with BRITS. The publicly available implementation of $E^{2}$ is unfortunately incompatible with our current setup, precluded the possibility of directly comparing $E^{2}$ with other models in our evaluation.
@@ 5.3 Experimental Results
The consolidated analysis of both imputation and classification tasks, as reflected in the experimental results, underscores the strengths and areas for improvement of the Conditional Self-Attention Imputation (CSAI) model across various healthcare datasets.
@@
Table 2 reflects CSAI’s robust classification capabilities, with competitive Area under the ROC Curve (AUC) scores across various masking ratios. It demonstrates the Conditional Self-Attention Imputation (CSAI) model’s overall effectiveness, with notable performance in the eICU, PhysioNet, and MIMIC_59f datasets. This superior performance, especially notable in the PhysioNe dataset, CSAI’s effectiveness in handling the complex task of predicting outcomes based on imputed data. However, in the MIMIC_89f dataset, CSAI is slightly underperformed compared to BRITS-GRU. It not only indicates the robustness of the BRITS architecture and underscores the strength of the BRITS framework; but also validates its selection as the foundational structure for the CSAI model. The slightly lower performance in the MIMIC_89f dataset could be due to various factors, such as the inherent complexity of the dataset or specific nuances in missing data patterns that might require additional model tuning.
Our comprehensive experimental evaluation of the Conditional Self-Attention Imputation (CSAI) model across diverse healthcare datasets demonstrates its effectiveness in addressing both imputation and classification tasks. CSAI’s adaptability to the complexities inherent in healthcare data, evidenced by its consistent performance across varying masking ratios, highlights its resilience to data sparsity. This is a significant advantage in real-world healthcare scenarios where incomplete data is a common challenge. The integration of domain-specific knowledge and the implementation of sophisticated imputation strategies in CSAI significantly contribute to its robustness, emphasizing the necessity of tailored approaches in the realm of healthcare analytics.
Despite the overall strong performance of CSAI, there are areas where CSAI’s performance indicates room for improvement. In the MIMIC_89f dataset, CSAI slightly lags behind BRITS-GRU, suggesting that further model refinement is needed. This disparity opens up avenues for future research, particularly in enhancing CSAI’s domain knowledge integration mechanisms and exploring the potential benefits of additional training data. Such refinements could improve CSAI’s robustness and effectiveness across various healthcare datasets, each with its unique characteristics and challenges.
@@
BVRIN 0.8876632792674032(0) 0.8842207935257471(0) 0.8846083295889355(0) 0.8327985224797132(0) 0.8331225972961287(0) 0.8273344786621177(0)
0.8853455913665667(10) 0.881934198142275(10) 0.8810832005813483(10) 0.8330776522899148(10) 0.8301506368159887(10) 0.8245590286971046(10)
BRITS 0.8866858256973604(0) 0.8852105575201445(0) 0.8856887012643835(0) 0.8282097938406451(0) 0.8277898698471445(0) 0.8241398505926232(0)
0.8858976145705622(10) 0.8854675509841966(10) 0.8859337578144677(10) 0.8254673610604841(10) 0.8255691464349476(10) 0.8183598901297295(10)
BRITS_GRU 0.8894446123008171(0) 0.8866442109252388(0) 0.8860531724480183(0) 0.8318754575867805(0) 0.830741815613582(0) 0.826876535038795(0)
0.8884356666666666(10) 0.8858544332092004(10) 0.8863321462371248(10) 0.8321182150682708(10) 0.8300872854458495(10) 0.8237612852706355(10)
@@ 5.4.2 Adjustment Factor Comparison
The experimental results depicted in the table indicate a nuanced relationship between the adjustment factor in the non-uniform masking strategy and the performance on imputation and classification tasks. An optimal adjustment factor demonstrated to be around 5, results in the lowest imputation error, suggesting that a balanced representation of features is crucial for accuracy. For classification, however, increasing the adjustment factor leads to higher errors and a marginal decline in AUC, highlighting that excessive weighting may not uniformly enhance performance across different machine-learning tasks. These findings reveal the delicate interplay between feature representation adjustments and task-specific model efficacy, underscoring the importance of calibration in the non-uniform masking approach for complex time series data.
@@ 7 Acknowledgments
8 Data availability
The datasets generated and/or analyzed during the current study are available from the corresponding author on reasonable request.
9 Ethics declarations
9.1 Conflict of interest
The authors declare no conflict of interest regarding the publication of this paper.